OpenStudy (anonymous):

Math Analysis: Find the sum of the given geometric series 12-6+3-3/2+...

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large a{ 1 - r^n \over 1 - r} }\) The ratio is -1/2 as we multiply by -1/2 all the time to get the next term in the sequence. a would be the first terms i.e 12. n would be the number of terms.

Parth (parthkohli):

Is this an infinite sequence?

OpenStudy (anonymous):

idk.. i haven't really learned this in class, our teacher just gave it to us for our final which is gunna start in 2 hrs..

OpenStudy (anonymous):

it says the answer is 8 but idk how they got that answer

Parth (parthkohli):

Honestly I have no idea too

OpenStudy (anonymous):

Apply the formula: S = a_1 / ( 1 -r) = 12 / 1.5 = 8

OpenStudy (anonymous):

why do we divide 12 by 1.5?

OpenStudy (anonymous):

where did the 1.5 come from

OpenStudy (anonymous):

Can you use this formula and @ParthKohli's instruction:

OpenStudy (anonymous):

can u explain it to me.. cause he said he had no clue what he was doing

Parth (parthkohli):

There is supposed to be a 1 - r^n....I don't know what n is supposed to be :/

OpenStudy (anonymous):

We know, it just because you confused us, and keep closing your post !?!

OpenStudy (anonymous):

so what would be n

OpenStudy (anonymous):

what was r in the first place

OpenStudy (anonymous):

S = a_1 / ( 1 -r) a_1 = 12 r = -1/2

OpenStudy (anonymous):

that made u get 1.5 from "a/1-r"

OpenStudy (anonymous):

whats the _ between 1 and a mean?

OpenStudy (anonymous):

im still confused

OpenStudy (anonymous):

Work out here, plug in the number, so I know what make you confused!

OpenStudy (anonymous):

ok so the formula is S=a1/1-r?

OpenStudy (anonymous):

a1=12 and to get r i would use that formula up there? "S=a1/1-r"

OpenStudy (anonymous):

I mean ratio

OpenStudy (anonymous):

which is 6?

OpenStudy (anonymous):

r=1.5? but how

OpenStudy (anonymous):

.5

OpenStudy (anonymous):

so r=.5

OpenStudy (anonymous):

sorry i mean -.5

OpenStudy (anonymous):

then why do we divide by 1.5?

OpenStudy (anonymous):

if r=-.5

OpenStudy (anonymous):

1 - ( -.5 ) = ....?

OpenStudy (anonymous):

k, i got it

OpenStudy (anonymous):

All you need is the formula!

OpenStudy (anonymous):

the formula for r=a2/a1?

OpenStudy (anonymous):

Your post is so confusing, make sure post the answer, so we can manage to help you out. Pls don't keep closing your post, jumping all around !!!

OpenStudy (anonymous):

wait so the formula for r=a2/a1

OpenStudy (anonymous):

You don't read at all, do you ??? S = a_1 / ( 1 -r) a_1 = 12 r = -1/2

OpenStudy (anonymous):

-1 would be -6 2 would be 12

OpenStudy (anonymous):

im sorry, i'm really tired rite now.. didn't get any sleep :\

OpenStudy (anonymous):

im sorry it's just hard atm when you don't sleep

OpenStudy (anonymous):

Because we already try so much to help you out, however you are the one neglect it !!!

OpenStudy (anonymous):

i just thought r=a2/a1 because we divided -6/12 to get our answer which was .5

OpenStudy (anonymous):

element 2/element 1

OpenStudy (anonymous):

r = any ratio between the next 2 numbers!

OpenStudy (anonymous):

what ifi replaced -6 with 24?

OpenStudy (anonymous):

-24? would i divide -24/12?

OpenStudy (anonymous):

or would it still be the same .5

OpenStudy (anonymous):

-.5

OpenStudy (anonymous):

12-6+3-3/2+.. But in this series the first one is 12, you can't break its rule!

OpenStudy (anonymous):

i know what i'm trying to say is what if the equation was like this 12-24+3-3/2+... would r still equal -.5?

OpenStudy (anonymous):

Yes, all the ratio should be equal, that's from geometric series definition!

OpenStudy (anonymous):

so it doesn't have to do with the 2ndnumber being multiplied by the 1st but the ratio between the two numbers

OpenStudy (anonymous):

*being divided*

OpenStudy (anonymous):

It's just all the basic from the formula! I'll leave it for you to learn for your own long run benefits!

OpenStudy (anonymous):

what if those two number were like in decimals..?

OpenStudy (anonymous):

like 12.5 and 7.5 how would i get the ratio from there?